# A Similarity Measure for Comparing Conversational Dynamics

### v1

@@ Abstract
In this work, we introduce a similarity measure for comparing conversations with respect to their dynamics. We design a validation procedure for testing the robustness of the metric in capturing differences in conversation dynamics and for assessing its sensitivity to the topic of the conversations. Finally, to illustrate the measure’s utility, we use it to analyze conversational dynamics in a large online community, bringing new insights into the role of situational power in conversations.
@@ A Similarity Measure for Comparing Conversational Dynamics
Sang Min Jungthanks:  Equal contribution. Cornell University sj597@cornell.edu          Kaixiang Zhang11footnotemark: 1 Cornell University kz88@cornell.edu
Cristian Danescu-Niculescu-Mizil Cornell University cristian@cs.cornell.edu
@@ 1 Introduction
In a conversation, individual utterances combine to form interactional patterns, such as exchange structures (e.g., brief exchanges vs. extended back-and-forth), changes in tone (e.g., from passive-aggressive to defusing), and conversational strategies (e.g., concessions or challenges). Each of these patterns contributes to shaping the conversation’s overall dynamics, and none of them alone is sufficient to characterize it Hua et al. (2024).
@@
However, developing a method for comparing conversations with respect to their overall dynamics presents several challenges. The first challenge is finding an appropriate way of representing the dynamics of a conversation: it is not sufficient to detect individual patterns separately, as done by prior work (e.g., speech acts, empathy, politeness, sarcasm) Ghosh et al. (2017); Oraby et al. (2017); Chhaya et al. (2018); Danescu-Niculescu-Mizil et al. (2013). Instead, a representation of the overall dynamics must capture how relevant patterns of different types connect to each other. For example, a passive-aggressive tone changing into a defusing tone leads to a very different dynamic than when a defusing tone is followed by a passive-aggressive tone.
@@
To validate the effectiveness of ConDynS and compare it with baseline measures using other representations or alignment methods, we introduce a human-in-the-loop procedure for generating labeled data. ConDynS recovers these labels with over 90% accuracy, substantially outperforming the baselines, while being robust against changes in topic.
Finally, we demonstrate how a similarity measure for conversational dynamics can enable new types of analysis by applying ConDynS to conversations from a large online debate community. First, it allows us to adapt standard similarity-based techniques—clustering, inter-group similarity, and intra-group diversity—to study conversational dynamics. Second, we use our measure to investigate which participants are more likely to influence the dynamics of a conversation, providing new insights into the role of situational power in conversations.
In summary, in this work we introduce a similarity measure for comparing conversational dynamics. We introduce a validation procedure that allows it to compare it with other baseline measures. And we showcase its use in analyzing online discussions. To encourage further applications, we also make our code publicly available as part of ConvoKit, including demos in other domains. [^1] https://convokit.cornell.edu
@@
2 Further Related Work
Evaluating conversation-level similarity. Prior work on measuring conversation-level, rather than utterance-level, similarity is limited. Lavi et al. (2021) adapts edit distance to measure similarity of “dialogue flow”, by defining substitution cost based on the semantic similarity of utterances. Other methods Bhaumik et al. (2023) additionally consider semantic features specific to task-oriented interactions, such as agent intent. In contrast, ConDynS is not concerned with the topic or semantics of what is discussed, focusing solely on interactional patterns and the emerging dynamics.
Other measures focus on a few predefined features, such as dialog acts Enayet and Sukthankar (2022), sentiment Xu et al. (2019), and number of words per turn Appel et al. (2018). Our measure instead captures how multiple types of interactional patterns combine, without predefining what types of dynamics to consider.
Representing conversations as sequences. Even though not explicitly defining a similarity measure, several studies represent conversations as sequences of predefined dialog units and aggregate them to compare groups of conversations. Mirzakhmedova et al. (2023) introduces sequences of argumentation strategies previously defined by Morio et al. (2019). Wang and Cardie (2014) represents a conversation as the flow of sentiment in its utterances. Unlike the SoP representation we use, these sequence representations focus on a single predefined dimension of the conversation.
Synthetic conversations. LLMs have been used to generate and annotate datasets across NLP tasks (see Tan et al. (2024) for a survey) including in the conversational domain Wang et al. (2024); Louie et al. (2024); Liu et al. (2024), sometimes with expert human input Louie et al. (2024). We build on this work to design our validation procedure, which uses human-written summaries to generate conversations with labels for relative similarity.
@@ 3 Measure
Measuring the similarity between the dynamics of two conversations involves (1) representing these dynamics and (2) comparing them. Below, we discuss several options for these steps, which combine to form ConDynS and several baseline measures. Here we describe the general approach, and defer to Section 4 for the operationalization in the specific domain of online debate discussions.
@@ 3.1 Representing conversational dynamics
Conversational dynamics are complex, emerging from the progression and juxtapositions of multiple interaction patterns. Therefore, their representation must go beyond describing individual patterns separately (e.g., how polite each reply is, whether it is sarcastic or not, etc.), and instead capture how relevant patterns combine to form the conversation’s dynamics. Given that patterns emerge at multiple scales and are often overlapping, there is an inherent tradeoff between precisely representing a coherent progression and capturing all patterns.
@@ 3.2 Comparing dynamics
A straightforward approach to compare dynamics is to measure how well the interaction patterns in one conversation match with the ones in the other conversation. Our approach additionally recognizes the role of the order in which patterns appear and quantifies how well the patterns in the two conversations are aligned.
@@
  • Naive prompting: We prompt a language model to give a similarity score between 0 and 100100100 in terms of their dynamics. The prompt is included in Figure 11 and 12.
@@
where $s_{i}\in[0,1]$ indicates how much $p^{i}_{c_{1}}\in P_{c_{1}}$ contributes to the alignment with the dynamics of $c_{2}$. In addition to rewarding ones that also appear in $c_{2}$, the score is designed to penalize patterns that: (1) appear out of order in $c_{2}$, and (2) are separated in $c_{2}$ from the previous pattern in the $c_{1}$ sequence (e.g., by other patterns that only appear in $c_{2}$). At the extremes, a pattern $p^{i}_{c_{1}}$ that does not appear in $c_{2}$ will receive a score $s_{i}=0$ and a pattern $p^{i}_{c_{1}}$ that also appears in $c_{2}$ immediately after a pattern matching $p^{i-1}_{c_{1}}$ will have a score $s_{i}=1$.
@@ 4 Data and Operationalization
Online debate discussions. To validate and demonstrate applications of ConDynS, we use a dataset of conversations from the ChangeMyView subreddit (CMV), retrieved from ConvoKit Chang et al. (2020). The objective of this platform is for participants (Challengers) to persuade the original poster (OP) to change their viewpoint on an opinion they hold. [^3] We follow prior work and consider a conversation to be one linear reply-chain starting with the first comment to the original post introducing the to-be-changed opinion Chang and Danescu-Niculescu-Mizil (2019); Hua et al. (2024). The dataset includes all conversations in the subreddit from its inception, in 2015, up to 2018, and is thus not polluted by content generated by large language models.
@@
This setting has several properties that make it particularly suitable for developing a similarity metric for conversational dynamics. First, it has been a resource for many studies analyzing how conversational features connect to different outcomes—such as successful persuasion Tan et al. (2016); Priniski and Horne (2018); Monti et al. (2022); Wei et al. (2016) or conversation derailment Altarawneh et al. (2023); Kementchedjhieva and Søgaard (2021); Chang and Danescu-Niculescu-Mizil (2019)—documenting its richness in conversational dynamics. Second, a key feature of the dataset is the “delta” ($\Delta$) mechanism through which the OP can award a $\Delta$ to a Challenger that successfully changed their view. This mechanism provides explicit persuasion labels for each conversation, which we will use to interpret our results. Finally, Hua et al. (2024) developed the SCDs procedure on this dataset. As such, they distribute human-written SCDs and provide a validated procedural prompt for automatically generating SCDs, which allows grounding our method and validation procedure in an established framework.
@@
Operationalization. To apply ConDynS and baselines in the context of CMV, we make several implementation choices. The code we release is modular, making it easy to swap specific components to facilitate adaptation to other settings. We use Google’s Gemini 2.0 Flash model Anil et al. (2024) via their API for ConDynS, including generating SCDs, extracting SoP, and quantifying the alignment of dynamics. [^4] We also experimented with OpenAI’s latest chatgpt-4o model Achiam et al. (2024) on the validation set. The results for each measure compare similarly, and are included in Appendix D. We opt for Gemini because it is more cost-effective. To generate SCDs, we use the procedural prompt validated by (Hua et al., 2024). To measure the alignment scores $s_{i}$, we use a few-shot in-context learning prompt with human-constructed examples (showing scoring and reasoning) to quantify alignment. Prompts are included in Appendix A.
@@ 5 Validation
To generate the anchor-positive-negative triplets, we make use of the human-written SCDs provided by Hua et al. (2024) for a subset of 50 ChangeMyView conversations. These are paired on outcome, such that each conversation that derails into a personal attack is matched with a similar-topic, similar-length conversation that does not. [^7] According to (Hua et al., 2024), the SCDs were written by annotators based on truncated transcripts, such that they could not know the actual outcome of the conversations while writing the summaries. This data allows us to create 50 triplets with known relative similarity.
Sensitivity to topical context. Ideally, a reliable similarity metric for conversational dynamics would not be confounded by topic. To check which measure best embodies this ideal, we impose the topic of the simulated conversations in the triplet to obtain the following conditions: (1) both the positive and negative conversations are assigned the same topic as the anchor; (2) both positive and negative conversations are assigned a different topic from the anchor; and (3) an adversarial condition in which the positive counterpart has a different topic from the anchor, while the negative counterpart is assigned the same topic. The details of the operationalization, including the prompts for identifying and assigning topics are in Appendix C.
Validation results. Table 1 shows how accurately each similarity measure distinguished between similar (anchor-positive) and dissimilar (anchor-negative) pairs of conversations. ConDynS outperforms all baselines based on matching, in all topic conditions, highlighting the importance of accounting for the order of the interaction patterns through our alignment procedure. Furthermore, aligning the SoP to the transcript, and thus allowing for better recall of interaction patterns, results in additional gains over SoP-to-SoP alignment.
@@ 6 Applications
To ensure that results are not driven by basic structural differences like participant count or conversation length, we restrict all our analyses to dyadic conversations of 4-6 utterances, [^8] We select these conversations that are roughly 5 utterances long because they are long enough to allow dynamics to develop, while also being quite common. We obtain similar results without this strict length control (enforcing only that conversations have at least 4 utterances). which include only the OP and one Challenger (with the Challenger always starting the conversation). We leave it to future work to explore dynamics in group conversations.
@@ 6.1 Similarity-based data analysis
Clustering. To explore common dynamics in CMV, we cluster a random sample of 200200200 conversations from the last year of the data (201820182018) using hierarchical clustering with ConDynS. This involves computing the similarity between all possible pairs of conversations, for a total of 199001990019900 comparisons.
We qualitatively characterize the two top-level clusters by exploiting the natural language representation used by ConDynS. Specifically, we aggregate all patterns that receive an alignment score $s_{i}>0.5$ when measuring the similarity of two conversations in the same cluster. We compare the aggregated patterns from the two clusters using a Bayesian distinguishing-word analysis Monroe et al. (2008), and manually investigate the most distinguishing patterns. The results are summarized in Table 2 and examples of corresponding patterns are provided in Table 5.
@@
Cluster 2’s wide range of conversational strategies also suggests an argumentative or potentially contentious interaction. The majority of the speakers express disagreement with the other’s argument. The speakers ask a lot of rhetorical questions in their responses. They use straw man fallacies and philosophical arguments and have to clarify their reasoning often via examples and analogies. Conversations in Cluster 1, on the other hand, use detailed elaboration to help others understand their arguments. They are more likely to agree and acknowledge the validity of the other speaker’s points and concerns; if not, they will compromise and concede to points where they share perspectives.
Figure 3: Outline for applying ConDynS to different analyses supported by similarity measures. Statistical significant differences marked with ∗∗∗ ($p<0.001$).
@@
Table 2: Summary of qualitative analysis of the two identified clusters. Examples of corresponding patterns are included in Table 5.
@@
Inter-group similarity. We can further support this interpretation by comparing these automatically detected clusters with a set of conversations that are known to be persuasive. We sample a set of 100100100 conversations where the OP awarded a $\Delta$ (henceforth set $\Delta$), and a corresponding set of 100100100 corresponding conversations which were not awarded a $\Delta$ (henceforth set $\neg\Delta$). [^9] We follow Tan et al. (2016) and select conversations triggered by the same post to control for topic and OP. There is no overlap between these conversations and those in our random sample. We find that, as suggested by our qualitative analysis, conversations in Cluster 1 are more similar to those that are known to be persuasive (set $\Delta$) than to those that are not (set $\neg\Delta$): mean ConDynS of 0.390.390.39 vs. 0.290.290.29, $p<0.001$ per a Mann Whitney U-test. It is worth noting that this difference remains significant ($p<0.001$) even if we discard all conversations from Cluster 1 that received a $\Delta$, showing that our method can identify conversations that have persuasive-like dynamics even though their persuasiveness is not explicitly acknowledged by the OP. [^10] In contrast, Cluster 2’s similarity to the two labeled sets is not significantly different.
Intra-group diversity. Finally, we demonstrate the use of our measure to analyze the diversity of dynamics in a set of conversations by calculating intra-group similarity of set $\Delta$ and set $\neg\Delta$, respectively. Persuasive conversations are significantly more similar to each other than those in which the persuasive attempt fails (mean ConDynS 0.520.520.52 vs. 0.390.390.39, $p<0.001$ according to Mann Whitney U-test; distribution shown in Figure 16).
Figure 4: Similarity between two conversations in which a speaker has the role of OP vs. between two conversations in which the same speaker is the Challenger. The difference is statistically significant ($p<0.01$).

### v2

@@ Abstract
In this work, we introduce a similarity measure for comparing conversations with respect to their dynamics. We design a validation procedure for testing the robustness of the metric in capturing differences in conversation dynamics and for assessing its sensitivity to the topic of the conversations. To illustrate the measure’s utility, we use it to analyze conversational dynamics in a large online community, bringing new insights into the role of situational power in conversations.
@@ A Similarity Measure for Comparing Conversational Dynamics
Sang Min Jungthanks: Equal contribution. Cornell University sj597@cornell.edu          Kaixiang Zhang11footnotemark: 1 Cornell University kz88@cornell.edu          Cristian Danescu-Niculescu-Mizil Cornell University cristian@cs.cornell.edu
@@ 1 Introduction
In a conversation, individual utterances combine to form interactional patterns, such as changes in tone (e.g., from passive-aggressive to defusing), conversational strategies (e.g., analogies, concessions, or challenges), and interaction sequences (e.g., extended back-and-forth vs. one-sided rants). Each of these patterns contributes to shaping the conversation’s overall dynamics, but none of them alone is sufficient to characterize it Tannen (2005); Hua et al. (2024).
@@
However, developing a method for comparing conversations with respect to their overall dynamics presents several challenges. The first challenge is finding an appropriate way of representing the dynamics of a conversation: it is not sufficient to detect individual patterns separately (e.g., speech acts, empathy, politeness, sarcasm), as done by prior work Ghosh et al. (2017); Oraby et al. (2017); Chhaya et al. (2018); Danescu-Niculescu-Mizil et al. (2013). Instead, a representation of the overall dynamics must capture how relevant patterns of different types connect to each other. For example, a passive-aggressive tone changing into a defusing tone leads to a very different dynamic than when a defusing tone is followed by a passive-aggressive tone.
@@
To validate the effectiveness of ConDynS and compare it with baseline measures using other representations or alignment methods, we introduce a human-in-the-loop procedure for generating labeled data. ConDynS recovers these labels with over 90% accuracy, substantially outperforming the baselines, while being robust against topical confounds.
We further demonstrate how a similarity measure for conversational dynamics can enable new types of analysis by applying ConDynS to conversations from a large online debate community. First, it allows us to adapt standard similarity-based techniques—clustering, inter-group similarity, and intra-group diversity—to study conversational dynamics. Second, we use our measure to investigate which participants are more likely to influence the dynamics of a conversation, providing new insights into the role of situational power in conversations.
In summary, in this work we:
  • introduce a similarity measure for comparing conversational dynamics;
  • propose a validation procedure that enables comparison against baseline measures;
  • use our measure to provide new insights into the role of situational power in conversations.
We additionally explore the versatility of our measure by applying it to two other conversational domains, including scripted casual conversations between friends and non-English discussions held in a collaborative setting. To encourage further use and development, we release the code for ConDynS publicly as part of ConvoKit, including demos on multiple datasets. [^1] https://convokit.cornell.edu
@@
2 Background and Related Work
Conversational dynamics. We use the term “conversational dynamics” to denote how different interactional patterns combine and unfold over the course of a conversation to shape its overall temporal trajectory. This is arguably a broad and under-defined concept, as it depends on what type of interactional patterns are considered most salient in a given context, and how these patterns are identified and tracked over the course of the interaction. Prior literature has studied it by focusing on the temporal sequence of specific utterance-level patterns in isolation, such as argumentation strategies Morio et al. (2019); Mirzakhmedova et al. (2023), sentiment Wang and Cardie (2014), communicative acts Liao et al. (2023), or dialog acts Zhao et al. (2022). Other work has focused on tracking patterns that only emerge at a higher multi-turn or conversational-level, such as turn-taking Sacks et al. (1974), time-sharing Zhang et al. (2025), coordination Fusaroli and Tylén (2016), or changes in topical entropy Fischer and Ram (2024).
In this work, we follow Hua et al. (2024) to take a holistic approach that does not limit the types of patterns that can contribute to the overall dynamics. By adopting this flexible framework, our measure can compare dynamics involving patterns that are potentially novel or domain-specific, rather than imposing strict theoretical constraints in advance.
Measuring conversation-level similarity. Prior work on measuring conversation-level, rather than utterance-level, similarity is limited. Lavi et al. (2021) adapts edit distance to measure similarity of “dialogue flow”, by defining substitution cost based on the semantic similarity of utterances. Other methods Bhaumik et al. (2023) additionally consider semantic features specific to task-oriented interactions, such as agent intent. In contrast, ConDynS is not concerned with the topic or semantics of what is discussed, focusing solely on interactional patterns and the emerging dynamics.
Other measures focus on a few predefined features, such as dialog acts Enayet and Sukthankar (2022); Zhao et al. (2022), sentiment Xu et al. (2019), and number of words per turn Appel et al. (2018). Our measure instead compares dynamics that emerge from how multiple types of interactional patterns combine and unfold.
Synthetic conversations. LLMs have been used to generate and annotate datasets across NLP tasks (see Tan et al. (2024) for a survey) including in the conversational domain Wang et al. (2024); Louie et al. (2024); Liu et al. (2024), sometimes with expert human input Louie et al. (2024). We build on this work to design our validation procedure, which uses human-written summaries to generate conversations with labels for relative similarity. We use real conversations for the rest of our analysis.
@@ 3 Measure
Measuring the similarity between the dynamics of two conversations involves (1) representing these dynamics and (2) comparing them. Below, we discuss several options for these steps, which combine to form ConDynS and several baseline measures. Here we describe the general approach, and defer to Section 4 for details about the operationalization in the specific domains we analyze in this work.
@@ 3.1 Representing conversational dynamics
Conversational dynamics are complex, emerging from the progression and juxtapositions of multiple interaction patterns. Therefore, their representation must go beyond describing individual patterns separately (e.g., how polite each reply is, whether it is sarcastic or not, etc.), and instead capture how relevant patterns combine to form the conversation’s dynamics. Given that patterns are often overlapping and can emerge at multiple scales—some being confined to a single utterance while others are spanning multiple utterances—there is an inherent tradeoff between precisely representing a coherent progression and capturing all patterns present in a conversation.
@@ 3.2 Comparing dynamics
A straightforward approach to compare dynamics would be to measure how well the interaction patterns in one conversation match with the ones in the other conversation. Our approach additionally recognizes the role of the order in which patterns appear and quantifies how well the patterns in the two conversations are aligned.
@@
  • Naive prompting: We prompt a large language model to compare two conversations in terms of their dynamics and give a similarity score between 11 and $100$. The prompt is in the Appendix in Figures 13 and 14.
@@
where $s_{i}\in[0,1]$ indicates how much $p^{i}_{c_{1}}\in P_{c_{1}}$ contributes to the alignment with the dynamics of $c_{2}$. In addition to rewarding patterns that also appear in $c_{2}$, the score is designed to penalize patterns that: (1) appear out of order in $c_{2}$, and (2) are separated in $c_{2}$ from the previous pattern in the $c_{1}$ sequence (e.g., by other patterns that only appear in $c_{2}$). At the extremes, a pattern $p^{i}_{c_{1}}$ that does not appear in $c_{2}$ will receive a score $s_{i}=0$ and a pattern $p^{i}_{c_{1}}$ that also appears in $c_{2}$ immediately after a pattern matching $p^{i-1}_{c_{1}}$ will have a score $s_{i}=1$.
@@ 4 Data and Operationalization
Online debate discussions. To validate and demonstrate applications of ConDynS, we use a dataset of conversations from the ChangeMyView subreddit (CMV), retrieved from ConvoKit Chang et al. (2020). The objective of this platform is for participants (Challengers) to persuade the original poster (OP) to change their viewpoint on an opinion they hold. [^3] We follow prior work and consider a conversation to be one linear reply-chain starting with the first comment to the original post introducing the to-be-changed opinion Chang and Danescu-Niculescu-Mizil (2019); Hua et al. (2024). The dataset includes conversations from the subreddit’s inception in 2015, up to 2018, and is thus not polluted by content generated by large language models. In this paper we use a total of 9,138 CMV conversations, selected as described in Sections 5 and 6.
@@
This setting has several properties that make it particularly suitable for developing a similarity metric for conversational dynamics. First, it has been a resource for many studies analyzing how conversational features connect to different outcomes—such as successful persuasion Tan et al. (2016); Priniski and Horne (2018); Monti et al. (2022); Wei et al. (2016) or conversation derailment Altarawneh et al. (2023); Kementchedjhieva and Søgaard (2021); Chang and Danescu-Niculescu-Mizil (2019)—documenting its richness in conversational dynamics. Second, a key feature of the dataset is the “delta” ($\Delta$) mechanism through which the OP can award a $\Delta$ to a Challenger that successfully changed their view. This mechanism provides explicit persuasion labels for each conversation, which we will use to interpret our results. Finally, Hua et al. (2024) developed the SCDs procedure on this dataset. As such, they distribute human-written SCDs and provide a validated procedural prompt for automatically generating SCDs, which grounds our method and validation procedure in an established framework.
Other conversation settings. As discussed in Section 2, we adopt a flexible framework for conversational dynamics in order to allow adapting the measure to other domains where different patterns might be at play. To explore the versatility of ConDynS, we apply it to two additional settings (Section 7). The first setting involves 50 fictional conversations from the Friends TV show Chen and Choi (2016) which are scripted to resemble everyday face-to-face interactions and reflect entertainment-driven dialogue. The second is a collaborative setting in a non-English language, specifically 100 conversations from the German Wikipedia talk-pages Hua et al. (2018).
@@
Operationalization. We release a modular implementation of ConDynS in ConvoKit, making it easy to swap specific components to facilitate adaptation to different settings. We use Google’s Gemini 2.0 Flash model’s API Anil et al. (2024) for generating SCDs, extracting SoP, and quantifying the alignment of dynamics. [^4] We also experimented with OpenAI’s chatgpt-4o model Achiam et al. (2024) on the validation set, without noting substantial changes in performance (Appendix D). To generate SCDs, we use the procedural prompt validated by Hua et al. (2024) for our main CMV setting, and modified versions that include domain-specific examples for the additional settings. All generated SCDs are distributed together with the respective datasets in ConvoKit. To measure the alignment scores $s_{i}$, we use a few-shot in-context learning prompt with human-constructed examples (showing scoring and reasoning) to quantify alignment. All prompts are included in Appendix A.
@@ 5 Validation
To generate the anchor-positive-negative triplets, we make use of the human-written SCDs provided by Hua et al. (2024) for a subset of 50 ChangeMyView conversations. These are paired on outcome, such that each conversation that derails into a personal attack is matched with a similar-topic, similar-length conversation that does not. [^7] According to Hua et al. (2024), the SCDs were written by annotators based on truncated transcripts, such that they could not know the actual outcome of the conversations while writing the summaries. This data allows us to create 50 triplets with known relative similarity.
Sensitivity to topical context. Ideally, a reliable similarity metric for conversational dynamics would not be confounded by topic. To check which measure best embodies this ideal, we control the topic of the simulated conversations in the triplet to obtain the following conditions: (1) both the positive and negative conversations are assigned the same topic as the anchor; (2) both positive and negative conversations are assigned a different topic from the anchor; and (3) an adversarial condition in which the positive counterpart has a different topic from the anchor, while the negative counterpart is assigned the same topic. The details of the operationalization, including the prompts for identifying and assigning topics are in Appendix C.
Validation results. Table 1 shows how accurately each similarity measure distinguished between similar (anchor-positive) and dissimilar (anchor-negative) pairs of conversations. ConDynS outperforms all baselines based on matching, in all topic conditions, highlighting the importance of accounting for the order of the interaction patterns through our alignment procedure. Furthermore, aligning the SoP to the transcript—and thus allowing for better recall of interaction patterns—results in additional gains over SoP-to-SoP alignment.
@@ 6 Applications
To ensure that results are not driven by basic structural differences like participant count or conversation length, we focus on conversations that involve only the OP and one Challenger (who always initiates) and that are at least 4 utterances long (and thus are long enough to allow dynamics to develop). We also consider a stricter length control, in which all conversations are between 4 and 6 utterances. Both conditions lead to similar qualitative and numerical results, with small changes in significance levels. In what follows we report the results with the strict length control, and report those without strict length control in Appendix F.
@@ 6.1 Similarity-based data analysis
Clustering. To explore common dynamics in CMV, we cluster a random sample of $200$ conversations from the last year of the data ($2018$) using hierarchical clustering with ConDynS. This involves computing the similarity between all possible pairs of conversations, for a total of $19,900$ comparisons.
We qualitatively characterize the two top-level clusters by exploiting the natural language representation used by ConDynS. Specifically, we aggregate all patterns that receive an alignment score $s_{i}>0.5$ when measuring the similarity of two conversations in the same cluster. We compare the aggregated patterns from the two clusters using a Bayesian distinguishing-word analysis Monroe et al. (2008), and manually investigate the most distinguishing patterns. The results are summarized in Table 2 and examples of corresponding patterns are provided in Tables 5 and 6 in the Appendix.
@@
Cluster 2’s wide range of conversational strategies also suggests an argumentative or potentially contentious interaction. The majority of the speakers express disagreement with the other’s argument. The speakers ask a lot of rhetorical questions in their responses. They use straw man fallacies and philosophical arguments and often have to clarify their reasoning via examples and analogies. Conversations in Cluster 1, on the other hand, use detailed elaboration to help others understand their arguments. They are more likely to agree and acknowledge the validity of the other speaker’s points and concerns; if not, they will compromise and concede to points where they share perspectives.
Figure 3: Outline for applying ConDynS to different analyses supported by similarity measures. Statistically significant differences marked with ∗∗∗ ($p<0.001$).
@@
Table 2: Summary of qualitative analysis of the two identified clusters. Examples of corresponding patterns are included in Tables 5 and 6 in the Appendix.
@@
Inter-group similarity. We can further support this interpretation by comparing these automatically detected clusters with a set of conversations that are known to be persuasive. We sample a set of $100$ conversations where the OP awarded a $\Delta$ (henceforth set $\Delta$), and a corresponding set of $100$ corresponding conversations which were not awarded a $\Delta$ (henceforth set $\neg\Delta$), while being triggered by same posts (thus controlling for topic and OP, following Tan et al. (2016)). There is no overlap between these sets and the random sample used for clustering.
We find that, as suggested by our qualitative analysis, conversations in Cluster 1 are more similar to those that are known to be persuasive (set $\Delta$) than to those that are not (set $\neg\Delta$): mean ConDynS of $0.39$ vs. $0.29$, $p<0.001$ per a Mann Whitney U-test. It is worth noting that this difference remains significant ($p<0.001$) even if we discard all conversations from Cluster 1 that received a $\Delta$, showing that our method can identify conversations that have persuasive-like dynamics even though their persuasiveness is not explicitly acknowledged by the OP. In contrast, Cluster 2’s similarity to the two labeled sets is not significantly different.
Intra-group diversity. Finally, we demonstrate the use of our measure to analyze the diversity of dynamics in a set of conversations by calculating intra-group similarity of set $\Delta$ and set $\neg\Delta$, respectively. Persuasive conversations are significantly more similar to each other than those in which the persuasive attempt fails (mean ConDynS $0.52$ vs. $0.39$, $p<0.001$ according to Mann Whitney U-test; distribution shown in Figure 20).
Figure 4: Similarity between two conversations in which a speaker has the role of OP vs. between two conversations in which the same speaker has the Challenger role. The difference is statistically significant ($p<0.01$).